• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. Howell

    Log entry time 08:19:24 on January13,2002

    Entry number 80225

    keyword=Shift Summary

    Summary:
    Most of the shift was used to determine the position of the targets. We had
    to assure that the beam will not hit the material of the target block when a cryo
    cell is put into beam position. With the first CW beam on the BeO target, we
    observed the beam spot to be at the extreme left edge of the foil (mostly
    hitting the Al frame when the BPMB abs(x) and abs(y) were less than 0.5 mm. The "left" side is as viewed looking from the dump to the target. Moving the
    beam to BPMB x = -4.0 mm and abs(y) < 0.5mm got the beam off the frame,
    causing the counting rate in the HRS scintillators to drop by a factor of about 9.
    Even with this adjustment, the beam spot is still about 7 mm from the center of
    the target in the horizontal direction and about 4 mm below the center vertically.

    Started taking runs with LH2 (with radiator) around 5:45 AM.

    Tasks accomplished:
    1. Measured beam position and profile on target.
    2. Determined condition of horizontal and vertical alignment of targets. Found
    that the target is 5 to 8 mm off center in all directions (x, y and z in target
    coordinate system).
    3. Started taking data with LH2 (with radiator) - this is #2 on the task list
    written on the board. Started with 1 uA beam current.

    Problems & Solutions:
    1. Found that the target alignment is pretty bad (off center by roughly 8 mm).
    * Solution:
    (a) request beam position BPMA: abs(x) < 1.0 mm, abs(y) < 0.5 mm
    BPMB: x = -4.0 mm, abs(y) < 0.5 mm
    (b) Maximum beam current = 5 uA



    Notes:
    1. 00:30 - Requested to close the hall (controlled access --> power permit).
    The LH2 target is now ready.
    2. 01:00 - Pulsed beam on BeO target for adjusting the beam tune parameters.
    Requested beam position in hall to be within 0.5 mm of origin on x and y
    of BPMA and BPMB.
    3. 02:00 - CW beam tuned to target, about 5 uA. Run 1468 accumulated while
    beam is being tuned. With beam spot within 1 mm on x of BPMA
    and BPMB, the beam hits the left side (as viewed looking back from the
    dump through the target) of the BeO target frame. This situation
    suggests that there is an inconsistency the target x position and the
    alignment of the BPMs.
    4. 03:00 - Stop run 1468 and replace BeO target with target 40-mil thick Aluminum target that has a square hole in center. Essentially no
    counts in HRS with this target, implying that the beam is not hitting
    the Al. Dave will step the target position to find the top edge of one
    target to verify the vertical offsets for the target. The HRS trigger
    rate will be used determine when the beam hits the frame.
    5. 03:45 - The target lifter stopped working. Dave requested an access into the
    Hall to reset target OIC.
    6. 05:45 - 15-cm long LH2 (with radiator) target put into beam position.
    7. 07:45 - Stopped run 1475. Turn the Big Box magnet current up to 500 A.
    There were two bad run on CODA (1476 and 1477). Started Run
    1477. The beam current on target is 1 uA, and the target is
    15-cm long LH2 (with radiator) target.